OpenStudy (anonymous):

Keith bought a hat and a poster for $26. Sarah’s family purchased four hats and two posters for $88. Write equations for each purchase. Solve your equations to find the cost of a hat and a poster.

OpenStudy (ganpat):

let hat = x and y = posters so, Keith = x + y = 26.. and Sarah, 4x + 2y = 88... can u solve them ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help me

OpenStudy (ganpat):

 x = 18.. y = 26 - 18 = 8